OpenStudy (anonymous):

2. Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions and identify what each term and factor represents in terms of the cell phone plans.

OpenStudy (anonymous):

@Floridagirl10, it seems you need to start with defining a variable which counts the number of months of contract duration. Let's call this variable simply 'm'. Then forget about the fact that you have two contracts but only look at the costs for a single contract. If you would sign up for such a contract and the company tells you to pay 10 USD now and after that 2 USD's per month, then you would pay the following for each additional month of your contract: 1st month (m=1) : 10+2 = 12 2nd month (m=2) : 10+2+2 = 14 3rd month (m=3) : 10+2+2+2 = 16 etc. You probably can see the logic behind how to arrive at the overall costs. In this example I've jsut selected a startfee of 10 USD, but if there's no figure given, you can replace this number by 'S' and forget about an actual value. Same applies to the monthly fee, which you can replace by F if you don't know a value. The overall formula for such a calculation then comes down to: C (for costs) = S + F*m. Obviously, two different providers will require you to create two formulae of the same shape, but with specific/different values for S and F. Let me know if you don't understand this approach.

OpenStudy (anonymous):

Thank you @NLcircle